OpenStudy (anonymous):

Ayee I'm having a problem answering questions involving trig and unknown angles... Simple one sin2X=SinX... From the known values the Answer is 60... however I'm not sure how to show that it's 60

OpenStudy (sburchette):

You can use the double angle formula to rewrite sin2x. The formula states\[\sin2x=2\sin x \cos x\] So now we can divide both sides by sinxcosx and we get \[1/\cos x =2\] becuase the sinx will cancel out the top sinx. If we cross multiply we get \[\cos x=1/2\]. Now we just have to take the inverse cos to get the value of x.\[x=\arccos (1/2)=60\]

OpenStudy (anonymous):

Thank you VERYYY Much :D